a father is 5 times as old as his son if the sum of their age is 60years,how old is the father
Answer by ikleyn(52800)

.
If the son is x years old, then the father is 5x years old,


and from the condition you have this equation


    x + 5x = 60.


So,


    6x = 60  ====>  x = 60/6 = 10.


ANSWER.  The son is 10 years old;  the father is 5*10 = 50 years old.